# YOLO-MARL: You Only LLM Once for Multi-agent Reinforcement Learning

### v1

@@ 4.3 Planning Function Generation
To be more concise, given any processed state $S_{I}$, we define an assignment planning function as $\mathcal{F_{\mathcal{T}}}(S_{I})\to\mathcal{T}_{i}\in\mathcal{T}$, where $\mathcal{T}=\{\mathcal{T}_{1},...,\mathcal{T}_{n}\}$ is a set of target assignments that each agent can take. We define the assignment set $\mathcal{T}$ over the action space such that an action can belong to multiple assignments and vice versa. For example, if the assignment space is defined as $\mathcal{T}=\{Landmark\_0,Landmark\_1\}$, and landmark 0 and landmark 1 are located at the top right and top left positions relative to the agent respectively, then taking the action ”UP” can be associated with both assignments. Conversely, we can have multiple actions correspond to an assignment. For instance, moving towards ”Landmark 0” may involve actions like ”UP” and ”RIGHT”. This is referred to the red module in Fig. 1(c) and more information of generated function refer to Appendix D.
4.4 MARL training with Planning function incorporation
@@
To incorporate the planning function into MARL training, we add an extra reward term to the original reward provided by environments. Specifically, we define the final reward $R$ used by the critic as: $R=r+\sum_{i}\Delta r_{i}$, where $r$ is the original reward from the environment and $\Delta r_{i}$ is an additional reward signal for each agent $i$. The term $\Delta r_{i}$ is based on whether the action taken by the agent’s actor aligns with the assigned task from the planning function. If the agent’s action belongs to the assigned task, it receives a reward $r^{\prime}$ as its $\Delta r_{i}$ and will receives a penalty $p^{\prime}$ as its $\Delta r_{i}$ if not.
@@
Notably, we do not need to interact with the LLM during the entire training process, nor do we need to call the planning function after the policy has been trained. The training process $MARL_{alg}(R)$ takes $R$ as the reward function, uses the same state and action space definitions and follows the standard MARL algorithms and evaluation metrics within the literature, such as Yu et al. (2022), Rashid et al. (2018), and Lowe et al. (2020). This makes our method highly efficient compared to approaches that interact with LLMs throughout the training process or use LLMs as agents as shown in the greed box in Fig.1(d). In practice, using the LLM’s API to generate the planning function incurs minimal cost—less than a dollar per environment—even when using the most advanced LLM APIs.
5 Experiments
In this section, we evaluate our method across three different environments: MPE, LBF, and SMAC. We use claude-3-5-sonnet-20240620 for the experiments.***We mainly use the Claude 3.5 Sonnet model for the LLM in our work: https://www.anthropic.com/news/claude-3-5-sonnet
5.1 Setup
@@
Baselines. In our experiments, we compare the MARL algorithm MADDPG (Lowe et al., 2020), MAPPO (Yu et al., 2022) and QMIX (Rashid et al., 2018) and set default hyper-parameters according to the well-tuned performance of human-written reward, and fix that in all experiments on this task to do MARL training. Experiment hyper parameters are listed in Appendix.
@@
Metrics. To assess the performance of our method, we use win rate as the evaluation metric on the SMAC environment, and the mean return in evaluation for all other environments. During evaluation, we rely solely on the default return values provided by the environments for both the baseline and our method, ensuring a fair comparison.
5.2 Results
@@
Table 1: Comparison between YOLO-MARL and MARL in the LBF environment across three seeds. The highest evaluation return means during training are highlighted in bold. The corresponding results can be found in Figure 2. The M means one million training steps. We run all the experiments on the same machine.
@@
Level-Based Foraging. Level-Based Foraging (LBF) (Papoudakis & Schäfer, 2021) is a challenging sparse reward environment designed for MARL training. In this environment, agents must learn to navigate a path and successfully collect food, with rewards only being given upon task completion. To evaluate our framework in a cooperative setting, we selected the 2-player, 2-food fully cooperative scenario. In this setting, all agents must work together and coordinate their actions to collect the food simultaneously. The environment offers an action space consisting of [NONE, NORTH, SOUTH, WEST, EAST, LOAD], and we define the task set as [NONE, Food i, …, LOAD]. Using the relative positions of agents and food items, we map assigned tasks to the corresponding actions in the action space and calculate the reward based on this alignment. We evaluated our framework over 3 different seeds, with the results shown in Figure 2 and Table 1. LLM assist the MARL algorithm by providing reward signals, our framework significantly outperformed the baseline, achieving a maximum improvement of 105 % in mean return and a 2x faster convergence rate among all tested MARL algorithms. According to the results, our framework is effective across all the baseline algorithms, with particularly large improvements observed in QMIX and MADDPG, and a faster convergence rate for MAPPO. To assess the variability in the quality of our generated functions, we present the results of three different generated functions in Figure 8 and Table 3 in Appendix B.1. The results demonstrate that our framework consistently generates high-quality functions, with each achieving similar improvements across all baseline algorithms.
@@
Multi-Agent Particle Environment. We evaluate our framework in Multi-Agent Particle Environment (MPE) (Lowe et al., 2020) simple spread environment which is a fully cooperative game. This environment has N agents, N landmarks. At a high level, agents must learn to cover all the landmarks while avoiding collisions. It’s action space is consist of [no_action, move_left, move_right, move_down, move_up]. We define the assignment for each agent to take to be [Landmark_i,…,No action]. During training, based on the global observation, we obtain the relative position of each agent with respect to the landmarks. Similar to LBF, we map each assignment of agent back to the corresponding action space and then reward the action of policy in action space level. We evaluate our approach on 3-agent and 4-agent scenarios using QMIX and MADDPG as baselines. As shown in Figure 3, our framework(colored line) outperform the baseline(black line) algorithm in mean returns by 7.66% and 8.8% for 3-agent scenario, and 2.4% and 18.09% for 4-agent scenario with QMIX and MADDPG respectively. These improvements demonstrate the effectiveness of our framework in enhancing coordination among agents to cover up all the landmarks.
(a) MADDPG
(b) QMIX
Figure 3: MPE simple spread scenario 3 agents results. The solid lines indicate the mean performance, and the shaded areas represent the range (minimum to maximum) across 3 different generated planning function.
(a) MADDPG
(b) QMIX
Figure 4: MPE simple spread scenario 4 agents results. The solid lines indicate the mean performance, and the shaded areas represent the range (minimum to maximum) across 3 different generated planning function.
StarCraft Multi-Agent Challenge environment. The StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al., 2019) simulates battle scenarios where a team of controlled agents must destroy an enemy team using fixed policies within a limited number of steps. We tested our method on three different maps: 3M, 2s vs 1sc, and 2c vs 64zg. The action space in the environment consists of [none, stop, move north, move south, move west, move east, attack enemy 1,…attack enemy n], where n is the total number of enemies on the map. This action space becomes increasingly complex depending on the number of enemies the agent has to engage, particularly in the 2c vs 64zg map, which contains 64 enemies and offers 70 possible actions.
In our experiments, we define the assignment space simply as [Move, Attack, Stop, None (for dead agents)]. We tested the performance of MAPPO, and the results for SMAC are shown in Figure 5. As indicated by the figure, even though we provide simple assignments that may be far from optimal instructions, our framework still achieves comparable results on certain maps. This demonstrates that our framework remains competitive, even in environments requiring strategic movements. We also explore the sparse reward case for this environment where the win rate of baseline algorithms is always closed to 0 while we generate a planning reward function pairs that outperform baseline. We suggest this pair generation as a potential future work and leave this discussion to the Sec 7.
(a) 3m
(b) 2s vs 1sc
(c) 2c vs 64zg
Figure 5: Results for 3 maps on SMAC environment: Average win rate comparison with our method for MAPPO baseline on 3 maps: 3m, 2s vs 1sc and 2c vs 64zg across 3 different seeds and the solid lines indicate the mean performance.
6 Ablation Study
In this section, we conduct the ablation studies mainly in LBF 2 players 2 food fully cooperative environment since rewards in LBF are sparser compared to MPE and SMAC (Papoudakis & Schäfer, 2021). We refer to 1 for more information about the environment. Due to page limitation, we also leave some discussions and figures in Appendix B.
6.1 Comparison between YOLO-MARL with and without Strategy Generation
@@
6.2 Comparison between YOLO-MARL with and without State Interpretation
@@
As shown in Figure 10, the LLM is unable to infer the type of state and attempts to fetch environment information via a non-existent key if no preprocessing code provided. And if environment code is provided without dimensional context for each component, the LLM is likely to make random guesses. In both scenarios, the absence of explicit state interpretation hinders the LLM’s ability to generate accurate and executable planning functions. These failures underscore the importance of the State Interpretation Module in bridging the gap between vectorized observations and the LLM’s requirement for semantically meaningful input.
@@
6.3 Comparison between YOLO-MARL and reward generation
In this section, we compare our YOLO-MARL method with approaches that utilize the LLM for reward generation without reward function template. We explore two scenarios: reward generation without feedback and reward generation with feedback.
@@ 6.3 Comparison between YOLO-MARL and reward generation
To investigate whether iterative refinement could improve the LLM generated reward function, we supply the LLM with the generated reward function from the prior iteration and feedback on its performance. Despite this iterative process, the LLM still fails to output a suitable reward function for the LBF environment. The mean return of evaluations remains close to zero, as shown in figure 9. The generated reward functions for each iteration are provided in Appendix E.
7 Limitation and Future Work
We acknowledge that the performance of YOLO-MARL may be highly correlated with the LLM’s ability and we haven’t tested YOLO-MARL with other LLMs like GPT-o1 due to the tier5 user requirement, and there might be a gap of YOLO-MARL’s performance between the Claude-3.5 and GPT-o1.
For future work, we are enthusiastic about the potential for LLMs to further enhance MARL, particularly as their planning capabilities improve. Specifically, we envision combining reward generation with planning functions to boost the performance of existing MARL algorithms in fully sparse environments. In this approach, we prompt the LLM to generate both a planning function and a reward function that replaces the environment-provided reward, following the pipeline described in Section 4. The function-pair method may require further refinement, and we will explore it as a future direction. A preliminary test of this framework is provided in Appendix B.4.
8 Conclusion
@@
We propose YOLO-MARL, a novel framework that leverages the high-level planning capabilities of LLMs to enhance MARL policy training for cooperative games. By requiring only a one-time interaction with the LLM for each environment, YOLO-MARL significantly reduces computational overhead and mitigates instability issues associated with frequent LLM interactions during training. This approach not only outperforms traditional MARL algorithms but also operates independently of the LLM during execution, demonstrating strong generalization capabilities across various environments.

### v2

@@ IV-C Planning Function Generation
To be more concise, given any processed state $S_{I}$, we define an assignment planning function as $\mathcal{F_{\mathcal{T}}}(S_{I})\to\mathcal{T}_{i}\in\mathcal{T}$, where $\mathcal{T}=\{\mathcal{T}_{1},...,\mathcal{T}_{n}\}$ is a set of target assignments that each agent can take. We define the assignment set $\mathcal{T}$ over the action space such that an action can belong to multiple assignments and vice versa. For example, if the assignment space is defined as $\mathcal{T}=\{Landmark\_0,Landmark\_1\}$, and landmark 0 and landmark 1 are located at the top right and top left positions relative to the agent respectively, then taking the action “UP” can be associated with both assignments. Conversely, we can have multiple actions correspond to an assignment. For instance, moving towards “Landmark 0” may involve actions like “UP” and “RIGHT”.
The Planning Function Generation will only be required once for each new environment you try to use. After you interact with the LLM to get generated planning function, you can directly use it in the later training process with different MARL algorithms. This is referred to the red module in Fig. 1(c).
IV-D MARL training with Planning function incorporation
@@
To incorporate the planning function into MARL training, we add an extra reward term to the original reward provided by environments. Specifically, we define the final reward $R$ used by the critic as:
$R=r+\sum_{i}\Delta r_{i}.$ (1)
Here, $r$ is the original reward from the environment. For each agent $i$, $\Delta r_{i}$ is an additional reward or penalty that determined based on whether the action taken by the agent aligns with the task assigned by the planning function. Specifically:
$\Delta r_{i}=\begin{cases}r^{\prime},\text{if $a_{i}\in\mathcal{T}_{i}$}\\ p^{\prime},\text{if $a_{i}\notin\mathcal{T}_{i}$}\end{cases}$ (2)
@@
Notably, we don’t need to interact with the LLM during the entire training process, nor do we need to call the planning function after the policy has been trained. The training process $MARL_{alg}(R)$ takes $R$ as the reward function, uses the same state and action space. We follow the standard MARL algorithms and evaluation metrics within the literature, such as [8], [9], and [7]. Our method, as shown in the green box in Fig. 1(d), is highly efficient compared to approaches that interact with LLMs throughout the whole training process or directly use LLMs as agents. In practice, using the LLM’s API to generate the planning function incurs minimal cost—less than a dollar per environment—even when using the most advanced LLM APIs.
V Experiments
In this section, we evaluate our method across two different environments: MPE and LBF.
V-A Setup
@@
Baselines. In our experiments, we compare the MARL algorithm MADDPG[7], MAPPO[8] and QMIX[9] and set default hyper-parameters according to the well-tuned performance of human-written reward, and fix that in all experiments on this task to do MARL training. Experiment hyper parameters are listed in Appendix.
LLM As one of the most advanced LLMs available at the time of experimentation, Claude 3.5 Sonnet exhibited enhanced reasoning and code generation abilities critical for our framework’s strategy formulation and planning functions. This performance advantage led us to select claude-3-5-sonnet-20240620 as our primary LLM for all reported experiments. [^1] https://www.anthropic.com/news/claude-3-5-sonnet
@@
Metrics. To assess the performance of our method, we use the mean return in evaluation for all other environments. During evaluation, we rely solely on the default return values provided by the environments for both the baseline and our method, ensuring a fair comparison.
V-B Results
@@
TABLE I: Comparison between YOLO-MARL and MARL in the LBF environment across three seeds. The highest evaluation return means during training are highlighted in bold. The corresponding results can be found in Figure 2. The M means one million training steps. We run all the experiments on the same machine.
@@
Level-Based Foraging. LBF[1] is a challenging sparse reward environment designed for MARL training. In this environment, agents must learn to navigate a path and successfully collect food, with rewards only being given upon task completion. To evaluate our framework in a cooperative setting, we selected the 2-player, 2-food fully cooperative scenario. In this setting, all agents must work together and coordinate their actions to collect the food simultaneously. The environment offers an action space consisting of [NONE, NORTH, SOUTH, WEST, EAST, LOAD], and we define the task set as [NONE, Food i, …, LOAD]. Using the relative positions of agents and food items, we map assigned tasks to the corresponding actions in the action space and calculate the reward based on this alignment. We evaluated our framework over 3 different seeds, with the results shown in Figure 2 and Table I. LLM assist the MARL algorithm by providing reward signals, our framework significantly outperformed the baseline, achieving a maximum improvement of 105 % in mean return and a 2x faster convergence rate among all tested MARL algorithms. According to the results, our framework is effective across all the baseline algorithms, with particularly large improvements observed in QMIX and MADDPG, and a faster convergence rate for MAPPO. To assess the variability in the quality of our generated functions, we present the results of three different generated functions in Figure 6 and Table II in Appendix. The results demonstrate that our framework consistently generates high-quality functions, with each achieving similar improvements across all baseline algorithms.
@@
Multi-Agent Particle Environment. We evaluate our framework in MPE[7] simple spread environment which is a fully cooperative game. This environment has N agents, N landmarks. At a high level, agents must learn to cover all the landmarks while avoiding collisions. It’s action space consists of [no_action, move_left, move_right, move_down, move_up]. We define the assignment for each agent to take to be [Landmark_i,…,No action]. During training, based on the global observation, we obtain the relative position of each agent with respect to the landmarks. Similar to LBF, we map each assignment of agent back to the corresponding action space and then reward the action of policy in action space level. We evaluate our approach on 3-agent and 4-agent scenarios using QMIX and MADDPG as baselines. As shown in Figure 3, our framework(colored line) outperforms the baseline(black line) algorithm in mean returns by 7.66% and 8.8% for 3-agent scenario, and 2.4% and 18.09% for 4-agent scenario with QMIX and MADDPG respectively. These improvements demonstrate the effectiveness of our framework in enhancing coordination among agents to cover up all the landmarks.
(a) 3 Agents – MADDPG
(b) 3 Agents – QMIX
(c) 4 Agents – MADDPG
(d) 4 Agents – QMIX
Figure 3: Results for MPE simple spread environment with 3 agents (top row a and b) and 4 agents (bottom row c and d). The solid lines indicate the mean performance, and the shaded areas represent the range (minimum to maximum) across 3 different generated planning functions.
VI Ablation Study
In this section, we conduct the ablation studies mainly in LBF 2 players 2 food fully cooperative environment since rewards in LBF are sparser compared to MPE[1]. We refer to  V-B for more information about the environment.
VI-A Comparison between YOLO-MARL with and without Strategy Generation
@@
VI-B Comparison between YOLO-MARL with and without State Interpretation
@@
The LLM is unable to infer the type of state and attempts to fetch environment information via a non-existent key if no preprocessing code provided. And if environment code is provided without dimensional context for each component, the LLM is likely to make random guesses. In both scenarios, the absence of explicit state interpretation hinders the LLM’s ability to generate accurate and executable planning functions. These failures underscore the importance of the State Interpretation Module in bridging the gap between vectorized observations and the LLM’s requirement for semantically meaningful input.
@@
VI-C Comparison between YOLO-MARL and reward generation
In this section, we compare our YOLO-MARL method with approaches that utilize the LLM for reward generation without reward function template. We explore two scenarios: reward generation without feedback and reward generation with feedback. For the reward generation without feedback, the reward function is generated at the same stage as the planning function for fair comparison. This means that we generate the reward function before all the training process for each new environment. For the reward generation with feedback, we first generate a reward function just like the reward generation without feedback. And then, iteratively, we will run a whole training process on this environment and pass the feedback of this training performance to the LLM, combined with previous prompts and ask the LLM to refine the previous generated reward function.
@@ VI-C Comparison between YOLO-MARL and reward generation
To investigate whether iterative refinement could improve the LLM generated reward function, we supply the LLM with the generated reward function from the prior iteration and feedback on its performance. Despite this iterative process, the LLM still fails to output a suitable reward function for the LBF environment. The mean return of evaluations remains close to zero, as shown in figure 7.
VII Conclusion
@@
We propose YOLO-MARL, a novel framework that enhances MARL policy training by integrating LLMs’ high-level planning capabilities with a single interaction per environment. This design reduces computational overhead and mitigates instability issues associated with frequent LLM interactions during training. This approach not only outperforms traditional MARL algorithms but also operates independently of the LLM during execution, demonstrating strong generalization capabilities across various environments.